Find the present value of $700 due in the future under each of these conditions:

  1. 7% nominal rate, semiannual compounding, discounted back 10 years. Do not round intermediate calculations. Round your answer to the nearest cent.

  2. 7% nominal rate, quarterly compounding, discounted back 10 years. Do not round intermediate calculations. Round your answer to the nearest cent.

  3. 7% nominal rate, monthly compounding, discounted back 1 year. Do not round intermediate calculations. Round your answer to the nearest cent.

  4. Why do the differences in the PVs occur?
Answer #1

a.

Present Value = 700/(1 + 0.07/2)20

Present Value = $351.80

b.

Present Value = 700/(1 + 0.07/4)40

Present Value = $349.72

c.

Present Value = 700/(1 + 0.07/12)12

Present Value = $652.81